Miguel makes a boat display box for a school project. The display box is in the shape of a rectangular prism with a length of 29.2 cm, a height of 19.5 cm and a width of 10 cm.
Part A: round the given measurement of the display box to the nearest cm
Part B: what Is the exact volume of the display box? Show your work and label your answer.
Part C: Miguel centered the boat in the bottom of the display box. The boat has a length of 20 cm, a height of 12 cm, and a width of 6 cm. exactly how many cm are on either side of the boat once it is centered in the display box? Show your work.
I know how to help my son with A and B but I can't remember how to do C. Can you please help?

Answer by macston(5194)

.
V=volume;L=length=29.2 cm; W=width=10 cm; H=height=19.5cm
.
Part A:
Round to nearest centimeter:
L=29.2cm=29cm
W=10cm=10cm
H=19.5cm=20cm
.
Part B:
By asking for EXACT volume, the question implies
you should not use the rounded values.
V=L*W*H
V=(29.2cm)(10cm)(19.5cm)=5694 cubic centimeters.
.
Part C:
The box is 10 cm wide and the boat is 6 cm wide:
(10cm - 6cm)/2= 2cm left on each side
The box is 29.2 cm long and the boat is 20 cm long:
(29.2cm - 20 cm)/2= 4.6 cm left on each end.
Assuming the boat sits on the bottom of the box
(19.5cm-12cm)=7.5cm space above the boat
(3.75 cm over and under if the boat floats centered)